a person has $200 to spend on two goods X and Y whose respective prices are $7 and $9 draw the budget line? what happens to the original budget line if the budget falls by %50 ?what happen to the originale budget line if the price of X doubles? what happen to the originale budget line to be spent if the price of Y falls to $8?


1
Expert's answer
2021-10-29T11:03:23-0400

(a)

"P_x=7, G_x=\frac{200}{7}=28.57"

"P_y=9,G_y=\frac{200}{9}=22.22"



(b)

"200-(0.50\times 200)=100"

"P_x=7,G_x=\frac{100}{7}=14.29"

"P_y=9,G_y=\frac{100}{9}=11.11"



(c)

"P_x=14, G_x=\frac{200}{14}=14.29"




(d)

"P_y=8, G_y=\frac{200}{8}=25."